# Green’s matching: an efficient approach to parameter estimation in complex dynamic systems

### v1

@@ Abstract
Parameters of differential equations are essential to characterize intrinsic behaviors of dynamic systems. Numerous methods for estimating parameters in dynamic systems are computationally and/or statistically inadequate, especially for complex systems with general-order differential operators, such as motion dynamics. This article presents Green’s matching, a computationally tractable and statistically efficient two-step method, which only needs to approximate trajectories in dynamic systems but not their derivatives due to the inverse of differential operators by Green’s function. This yields a statistically optimal guarantee for parameter estimation in general-order equations, a feature not shared by existing methods, and provides an efficient framework for broad statistical inferences in complex dynamic systems.
@@ 1 Introduction
Dynamic systems have found broad applications in modeling the evolution of systems in various domains, such as the physical motion of objects (Blickhan, 1989), the evolutionary paths of genes (Wu et al., 2014; Chen et al., 2017), and the transmission dynamics of epidemics (Tian et al., 2021; Tan et al., 2023). Among these systems, their dynamic behaviours are usually characterized by differential equations:
@@
where $X_{i}(t)$ and $\boldsymbol{X}(t)=\left(X_{1}(t),\cdots,X_{p}(t)\right)^{T}$ are temporal trajectories of the target system, $\mathcal{D}_{i}^{K}$ is an order-$K$ differential operator, and $f_{i}(\cdot;\boldsymbol{\beta})$ is a driving function indexed by parameters $\boldsymbol{\beta}$. Typically, $\mathcal{D}_{i}^{K}$ and $f_{i}(\cdot;\boldsymbol{\beta})$ are two important ingredients that describe the intrinsic behaviors and interdependence among system variables. The former determines the derivatives of curves (e.g., velocity or acceleration) driven by the latter, and the form of $f_{i}(\cdot;\boldsymbol{\beta})$ is usually grounded in specific fields with $\boldsymbol{\beta}$ to be estimated from empirical data. Parameter estimation in dynamic systems facilitates the recognition of evolution patterns from observed data, thereby improving our understanding of system variables and enabling predictions of future dynamics.
@@
There is a vast literature on parameter estimation in dynamic systems, which can be broadly summarized into two frameworks. The first framework is known as “nonlinear least squares”, which estimates the parameters by fitting the forward solution of (1) to the observed data (Ramsay and Hooker, 2017). However, when $f_{i}(\cdot;\boldsymbol{\beta})$ in (1) is a nonlinear function, such a trajectory matching approach requires repeated equation solving in its estimation steps, leading to substantial computational demands for parameter estimation. Alternatively, the second framework utilizes smoothing approaches to approximate system trajectories and their derivatives for parameter estimation. This framework includes a family of collocation methods following a “first smoothing, then estimation” two-step strategy (Varah, 1982; Brunel, 2008; Gugushvili and Klaassen, 2012; Dattner and Klaassen, 2015), where the accuracy of the estimated parameters depends on the quality of the initial smoothing. To improve performances, iterative steps can be incorporated into collocation methods for a better approximation of curves (Ramsay et al., 2007; Niu et al., 2016; Yang et al., 2021). Nonetheless, these approaches may raise additional computational costs in their estimation steps. For many cases, especially for complex dynamic systems, the original two-step strategy receives more attention in statistical inferences (Wu et al., 2014; Chen et al., 2017; Zhang et al., 2020; Dai and Li, 2021) due to its simple implementation and efficient computation.
@@
Gradient matching (Brunel, 2008; Gugushvili and Klaassen, 2012) and integral matching (Dattner and Klaassen, 2015; Ramsay and Hooker, 2017) are two important collocation methods, which are not only employed for parameter estimations but can also be extended to more complicated tasks, such as equation discovery (Brunton et al., 2016; Champion et al., 2019) or causal network inference (Wu et al., 2014; Chen et al., 2017; Dai and Li, 2021) for dynamic systems. In general, these collocation methods focus on model fitting by matching the unknown parameters with trajectories in some equations, where the trajectories are either fully observed or substituted by their smooth approximations from discrete noisy data. Such an equation-matching approach avoid time-consuming equation solving in parameter estimation, thereby providing significant computational benefits for many statistical tasks of dynamic systems. In addition to the computational advantages, the statistical properties of gradient matching and integral matching have also been discussed in the literature (Brunel, 2008; Gugushvili and Klaassen, 2012; Dattner and Klaassen, 2015). These studies demonstrated that both methods obtain the root-$n$ consistency for parameter estimation in first-order systems.
@@
Given the importance of these two methods, they mainly focus on systems with known first-order differential operators rather than cases with unknown higher-order differential operators. The latter is more general, covering a broad range of systems with general-order evolution patterns, such as motion dynamics. For these complex systems, the direct extension of gradient matching and integral matching may require pre-smoothing of curves’ derivatives. Nonetheless, pre-smoothing derivatives can be inaccurate, and even inappropriate, when the actual curves are not smooth (e.g., the acceleration of a motion trajectory could be discontinuous with a sudden change of force). This can lead to significant biases and poor theoretical behaviors for parameter estimation. Targeting general-order dynamic systems, estimating parameters and differential operators is essential but rarely discussed in the literature. We urgently need to develop a computationally tractable and statistically efficient approach served for these purposes.
@@
In this article, we adopt the “first smoothing, then estimation” two-step strategy for parameter estimation in general-order dynamic systems. Our study focuses on cases where the trajectories of dynamic systems are discretely and noisily observed, and we apply the local polynomial regression (Fan and Gijbels, 2018) to pre-smooth the trajectories from discrete noisy data. The proposed method is called Green’s matching, which utilizes Green’s functions of differential operators (Duffy, 2015) to incorporate pre-smoothing of trajectories into parameter estimations. As an important mathematical tool, Green’s function is a kernel function considered an inversion operation of a differential operator. It finds diverse applications ranging from approximating forward solutions of differential equations (Wahba, 1973; Duffy, 2015) to statistical inferences of differential-equation-based models (Heckman and Ramsay, 2000; Boullé et al., 2022; Stepaniants, 2023). However, these studies do not focus on estimating parameters and differential operators simultaneously using discretely and noisily observed trajectory data. Our proposed method fills this gap, providing an efficient approach for statistical inference in general-order dynamic systems.
@@
Overall, the superiors of Green’s matching lie in both statistical and computational aspects. Firstly, our method utilizes Green’s functions to invert the general-order differential operators in (1), thereby eliminating the need for pre-smoothing curves’ derivatives within parameter estimations. This appealing feature reduces the dependence on pre-smoothing for parameter estimation in general-order systems, compared to both gradient matching and integral matching. Furthermore, if $f_{i}(\cdot;\boldsymbol{\beta})$ in (1) satisfies a general separability condition, Green’s matching would degenerate to the conventional least squares fitting, greatly reducing the computational efforts required for general-order systems.
@@
Under mild conditions, we establish the root-$n$ consistency and the asymptotic normality for Green’s matching. Moreover, we extend the the ideas of gradient matching and integral matching to general-order dynamic systems and investigate their asymptotic properties. The estimation biases and convergence rates of these methods are both justified and discussed. As can be seen, the estimation bias would increase as more pre-smoothing is used within the parameter estimation. Furthermore, we reveal that the gradient matching and integral matching may not achieve the root-$n$ consistency as the differential order of dynamic systems increases. In contrast, Green’s matching always obtains this optimal rate for parameter estimations.
@@
The rest of this article is organized as follows. In Section 2, we first propose Green’s matching for parameter estimation in general-order dynamic systems. The statistical properties of Green’s matching and other two-step methods are then established in Section 3. Moreover, we conduct extensive simulation studies in Section 4 to investigate the estimation behaviors of the two-step approaches and the other competitive methods (Ramsay et al., 2007; Yang et al., 2021). Finally, we apply Green’s matching for equation discovery of a nonlinear pendulum in Section 5, and present conclusions and discussions in Section 6. The code, data, and proof of this article can be found at https://github.com/Tan-jianbin/Statistical-Inference-in-General-order-Dynamic-Systems.
@@ 2.1 Local polynomial regression
With some conditions and a suitably chosen bandwidth $h_{k}$, $\widehat{D^{k}X}_{i}(\cdot)$ serves as a consistent estimator of ${D^{k}X}_{i}(\cdot)$ as $n\rightarrow\infty$ (Fan and Gijbels, 1995), although its statistical convergence rate may slow down as $k$ increases (Stone, 1982). To illustrate this phenomenon, we present several examples of local polynomial regressions in Fig. 1. We observe that the estimated zero derivatives of curves accurately match their real shapes for different sample sizes $n$. However, the estimation accuracy diminishes for the higher-order derivatives. This discrepancy is likely due to the difficulties in capturing high-order effects of curves from limited noisy data.
@@ 2.2 Equation matching for dynamic systems
where |||||$\cdot$||| | ⋅ | | is the Euclidean norm and $w(\cdot)$ is a known function on $[0,C]$ to assign weights for different matching equations. The fitting procedure above is called equation matching, which serves as a general framework to avoid equation solving for parameter estimation in dynamic systems. We here review two important equation-matching methods: gradient matching and integral matching.
@@
We call the equation matching based on $\mathcal{T}^{i,K}_{t,\boldsymbol{\beta},\boldsymbol{\omega}}(\cdot)$ or $\mathcal{T}^{i,K-1}_{t,\boldsymbol{\beta},\boldsymbol{\omega},\alpha_{i}}(\cdot)$ as the order $K$ or $(K-1)$ gradient matching, respectively, as both methods fit the parameters to the gradients of curves. Note that via the incorporation of integration, the order $(K-1)$ gradient matching reduces one differential order compared to the order $K$ gradient matching. Nonetheless, these approaches both involve pre-smoothing of high-order derivatives for general-order systems, thereby imposing additional pre-smoothing errors. This is even inappropriate when the smooth high-order derivative does not exist, as shown in Section 4.1, Therefore, we require a new approach capable of eliminating all high-order derivatives in matching equations for high-order systems.
@@ 2.3 Generalized integral matching by Green’s function
In this subsection, we propose an efficient approach called Green’s matching for parameter estimation, which removes the pre-smoothed derivatives via Green’s functions of differential operators (Duffy, 2015). In general, let $\mathcal{D}$ be a differential operator of functions on $[0,C]$. We define Green’s function of $\mathcal{D}$ as a kernel function $G(t,s)$ on $[0,C]^{2}$ satisfying $\mathcal{D}G(t,s)=\delta(t-s)$, where $\mathcal{D}$ is applied to the variable $t$, and $\delta(\cdot)$ is a Dirac delta function at zero. Given $G(t,s)$ of $\mathcal{D}$ and any functions $X(t)$ and $f(t)$ s.t. $\mathcal{D}X(t)=f(t)$, we have $X(t)=g(t)+\int_{0}^{C}G(t,s)f(s)\ \mathrm{d}s$, where $g(\cdot)$ is any function in the kernel space of $\mathcal{D}$, denoted as $\operatorname{Ker}(\mathcal{D}):=\left\{g(\cdot);\mathcal{D}g(t)=0,\forall t\in[0,C]\right\}$. In theory, $G(t,s)$ is considered an inverse operation for the operator $\mathcal{D}$, which would be used to reduce the differential order of the matching equations.
@@
where $\textbf{e}^{\mathbf{M}}$ is the matrix exponential of a matrix $\mathbf{M}$, $[\ ]_{m,m^{\prime}}$ extracts the $(m,m^{\prime})^{th}$ element of a matrix, $\boldsymbol{\omega}_{i}=(\omega_{i0},\cdots,\omega_{i(K-1)})^{T}$, and $\boldsymbol{A}(\boldsymbol{\omega}_{i})$ is a $K\times K$ matrix with its $K^{th}$ row being $\boldsymbol{\omega}_{i}^{T}$ and $[\boldsymbol{A}(\boldsymbol{\omega}_{i})]_{k,k^{\prime}}=\mathbb{I}(k=k^{\prime}-1)$ for $k\leq K-1$ and $k^{\prime}\leq K$, among $\mathbb{I}(\cdot)$ is an indicator function. Given $\boldsymbol{\psi}_{i}(t)$, $G_{i}^{K}(t,s)$ can be constructed as
@@
As a result, $\mathcal{T}^{i,\text{Gree}}_{t,\boldsymbol{\beta},\boldsymbol{\omega}_{i},\boldsymbol{\alpha}_{i}}(\boldsymbol{\hat{X}})=\mathcal{T}^{i,\text{Gree}^{\prime}}_{t,\boldsymbol{\beta},\boldsymbol{\omega}_{i},\tilde{\boldsymbol{\alpha}}_{i}}(\boldsymbol{\hat{X}})$ for all $\boldsymbol{\beta}$, $\boldsymbol{\omega}_{i}$, and $t$, implying that different choices of Green’s functions in (13) only result in a shift for the minimizer of $\boldsymbol{\alpha}$ and does not alter the minimizer of $\boldsymbol{\beta}$ and $\boldsymbol{\omega}$. Therefore, it is safe to employ the Green’s function constructed by (10) for Green’s matching.
@@ 2.4 Optimization scheme for Green’s matching
where $\boldsymbol{Z}_{i}=(\sqrt{w(t_{1})}\hat{X}_{i}(t_{1}),\cdots,\sqrt{w(t_{H})}\hat{X}_{i}(t_{H}))^{T}$, $\boldsymbol{\Psi}_{i}=(\sqrt{w(t_{1})}\boldsymbol{a}_{i}(t_{1}),\cdots,\sqrt{w(t_{H})}\boldsymbol{a}_{i}(t_{H}))^{T}$ with $\boldsymbol{a}_{i}(t)=(\{\boldsymbol{\psi}(t)\}^{T},-\int_{0}^{C}{G}_{K}(t,s)\hat{X}_{i}(s)\mathrm{d}s,\cdots,-\int_{0}^{C}{G}_{1}(t,s)\hat{X}_{i}(s)\mathrm{d}s)^{T}$, $\boldsymbol{\nu}_{i}=(\boldsymbol{\alpha}_{i}^{T},\boldsymbol{\omega}_{i}^{T})^{T}$, and $\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})=(\sqrt{w(t_{1})}\int_{0}^{C}f_{i}(\boldsymbol{\hat{X}}(s),s;\boldsymbol{\beta})G_{K}(t_{1},s)\mathrm{d}s,\cdots,\sqrt{w(t_{H})}$ $\int_{0}^{C}f_{i}(\boldsymbol{\hat{X}}(s),s;\boldsymbol{\beta})G_{K}(t_{H},s)\mathrm{d}s)^{T}$. In the following, we demonstrate procedures for solving (15), which summary is given in Algorithm 1.
@@
15:$\boldsymbol{\nu}_{i}=(\boldsymbol{\Psi}_{i}^{T}\boldsymbol{\Psi}_{i})^{-1}\boldsymbol{\Psi}_{i}^{T}\left\{\boldsymbol{Z}_{i}-\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})\right\}.$
@@
where $\boldsymbol{g}_{i}(\cdot)$ is a known function and $(\boldsymbol{\beta}_{1}^{T},\cdots,\boldsymbol{\beta}_{p}^{T})^{T}=\boldsymbol{\beta}$. This condition assumes that $f_{i}(\boldsymbol{X}(t),t;\boldsymbol{\beta})$ can be decomposed into a product of trajectory-related and parameter terms, a condition that is commonly adopted for statistical inference in dynamic systems (Brunton et al., 2016; Chen et al., 2017; Dai and Li, 2021). Under (16), $\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})$ in (15) satisfies $\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})=\boldsymbol{\Phi}_{i}\boldsymbol{\beta}_{i}$, where $\boldsymbol{\Phi}_{i}=(\sqrt{w(t_{1})}$ $\int_{0}^{C}\boldsymbol{g}_{i}(\boldsymbol{\hat{X}}(s),s)G_{K}(t_{1},s)\mathrm{d}s,\cdots,\sqrt{w(t_{H})}\int_{0}^{C}\boldsymbol{g}_{i}(\boldsymbol{\hat{X}}(s),s)G_{K}(t_{H},s)\mathrm{d}s)^{T}$. As a result, $\boldsymbol{\nu}_{i}$ and $\boldsymbol{\beta}_{i}$ can be estimated using the least squares fitting of (15) for each $i$. This approach can be applied to cases such as the gene regulatory network, oscillatory dynamic direction model, and the harmonic equations in Section 4, and also facilitates the equation discovery in Section 5.
For general cases, the minimizer of (15) may not have a closed form, and we instead employ iteration methods to solve the optimization problem. Letting $\boldsymbol{\nu}_{i}(\boldsymbol{\beta}):=(\boldsymbol{\Psi}_{i}^{T}\boldsymbol{\Psi}_{i})^{-1}\boldsymbol{\Psi}_{i}^{T}\left\{\boldsymbol{Z}_{i}-\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})\right\}$, the optimization (15) can be represented as

### v2

@@ Abstract
Parameters of differential equations are essential for characterizing the intrinsic behaviors of dynamic systems. Numerous methods for estimating parameters in dynamic systems are computationally and/or statistically inadequate, especially for complex systems with general-order differential operators, such as motion dynamics. This article presents Green’s matching, a computationally tractable and statistically efficient two-step method, which only requires approximating trajectories in dynamic systems but not their derivatives, due to the inversion of differential operators via Green’s function. This approach yields a statistically optimal guarantee for parameter estimation in general-order equations, a feature not shared by existing methods, and provides an efficient framework for broad statistical inference in complex dynamic systems.
@@ 1 Introduction
Dynamic systems have found broad applications in modeling the evolution of systems in various domains, such as the physical motion of objects (Blickhan, 1989), the evolutionary paths of genes (Wu et al., 2014; Chen et al., 2017), and the transmission dynamics of epidemics (Tian et al., 2021; Tan et al., 2023). Among these systems, their dynamic behaviors are usually characterized by differential equations:
@@
where $X_{i}(t)$ and $\boldsymbol{X}(t)=\left(X_{1}(t),\cdots,X_{p}(t)\right)^{T}$ are temporal trajectories of the target system, $\mathcal{D}_{i}^{K}$ is an order-$K$ differential operator, and $f_{i}(\cdot;\boldsymbol{\beta})$ is a driving function indexed by parameters $\boldsymbol{\beta}$. Typically, $\mathcal{D}_{i}^{K}$ and $f_{i}(\cdot;\boldsymbol{\beta})$ are two important ingredients that describe the intrinsic behaviors and interdependence among system variables. The former determines the derivatives of curves (e.g., velocity or acceleration) driven by the latter, and the form of $f_{i}(\cdot;\boldsymbol{\beta})$ is usually grounded in specific fields, with $\boldsymbol{\beta}$ to be estimated from empirical data. Parameter estimation in dynamic systems facilitates the recognition of evolution patterns from observed data, thereby improving our understanding of system variables and enabling predictions of future dynamics.
@@
There is a vast literature on parameter estimation in dynamic systems, which can be broadly summarized into two frameworks. The first framework is known as “nonlinear least squares,” which estimates the parameters by fitting the forward solution of (1) to the observed data (Ramsay and Hooker, 2017). However, when $f_{i}(\cdot;\boldsymbol{\beta})$ in (1) is a nonlinear function, such a trajectory-matching approach requires repeated equation solving in its estimation steps, leading to substantial computational demands for parameter estimation. Alternatively, the second framework utilizes smoothing approaches to approximate system trajectories and their derivatives for parameter estimation. This framework includes a family of collocation methods following a “first smoothing, then estimation” two-step strategy (Varah, 1982; Brunel, 2008; Gugushvili and Klaassen, 2012; Dattner and Klaassen, 2015), where the accuracy of the estimated parameters depends on the quality of the initial smoothing. To improve performance, iterative steps can be incorporated into collocation methods for a better approximation of curves (Ramsay et al., 2007; Niu et al., 2016; Yang et al., 2021). Nonetheless, these approaches may incur additional computational costs in their estimation steps. In many cases, especially for complex dynamic systems, the original two-step strategy receives more attention in statistical inference (Wu et al., 2014; Chen et al., 2017; Zhang et al., 2020; Dai and Li, 2021) due to its simple implementation and efficient computation.
@@
Gradient matching (Brunel, 2008; Gugushvili and Klaassen, 2012) and integral matching (Dattner and Klaassen, 2015; Ramsay and Hooker, 2017) are two important collocation methods, which are not only employed for parameter estimation but can also be extended to more complicated tasks, such as equation discovery (Brunton et al., 2016; Champion et al., 2019) or causal network inference (Wu et al., 2014; Chen et al., 2017; Dai and Li, 2021) for dynamic systems. In general, these collocation methods focus on model fitting by matching the unknown parameters with trajectories in equations, where the trajectories are either fully observed or substituted by smooth approximations from discrete noisy data. Such an equation-matching approach avoids time-consuming equation solving in parameter estimation, thereby providing significant computational benefits for many statistical tasks in dynamic systems. In addition to the computational advantages, the statistical properties of gradient matching and integral matching have also been discussed in the literature (Brunel, 2008; Gugushvili and Klaassen, 2012; Dattner and Klaassen, 2015). These studies demonstrated that both methods achieve root-$n$ consistency for parameter estimation in first-order systems.
@@
Despite their importance, these two methods mainly focus on systems with known first-order differential operators rather than cases with unknown higher-order differential operators. The latter setting is more general and covers a broad range of systems with general-order evolution patterns, such as motion dynamics. For these complex systems, direct extensions of gradient matching and integral matching may require pre-smoothing of curve derivatives. Nonetheless, pre-smoothing derivatives can be inaccurate, or even inappropriate, when the actual curves are not smooth (e.g., the acceleration of a motion trajectory could be discontinuous due to a sudden change of force). This can lead to significant biases and poor theoretical behavior for parameter estimation. Targeting general-order dynamic systems, estimating parameters and differential operators is essential but rarely discussed in the literature. There is an urgent need to develop a computationally tractable and statistically efficient approach for these purposes.
@@
In this article, we adopt the “first smoothing, then estimation” two-step strategy for parameter estimation in general-order dynamic systems. Our study focuses on cases where the trajectories of dynamic systems are discretely and noisily observed, and we apply local polynomial regression (Fan and Gijbels, 2018) to pre-smooth the trajectories from discrete noisy data. The proposed method is called Green’s matching, which utilizes Green’s functions of differential operators (Duffy, 2015) to incorporate pre-smoothing of trajectories into parameter estimation. As an important mathematical tool, Green’s function is a kernel function that is considered an inversion operation of a differential operator. It finds diverse applications ranging from approximating forward solutions of differential equations (Wahba, 1973; Duffy, 2015) to statistical inference for differential-equation-based models (Heckman and Ramsay, 2000; boullé2022learning; Stepaniants, 2023). However, these studies do not focus on estimating parameters and differential operators simultaneously using discretely and noisily observed trajectory data. Our proposed method fills this gap, providing an efficient approach for statistical inference in general-order systems.
@@
Overall, the superiority of Green’s matching lies in both statistical and computational aspects. First, our method utilizes Green’s functions to invert the general-order differential operators in (1), thereby eliminating the need for pre-smoothing curve derivatives within parameter estimation. This appealing feature reduces the dependence on pre-smoothing for parameter estimation in general-order systems, compared to both gradient matching and integral matching. Furthermore, if $f_{i}(\cdot;\boldsymbol{\beta})$ in (1) satisfies a general separability condition, Green’s matching degenerates to the conventional least squares fitting, greatly reducing the computational effort required for general-order systems.
@@
Under mild conditions, we establish root-$n$ consistency and asymptotic normality for Green’s matching. Moreover, we extend the ideas of gradient matching and integral matching to general-order dynamic systems and investigate their asymptotic properties. The estimation biases and convergence rates of these methods are both justified and discussed. As shown, the estimation bias increases as more pre-smoothing is used within parameter estimation. Furthermore, we reveal that gradient matching and integral matching may not achieve root-$n$ consistency as the differential order of dynamic systems increases. In contrast, Green’s matching always attains this optimal rate for parameter estimation.
@@
The rest of this article is organized as follows. In Section 2, we first propose Green’s matching for parameter estimation in general-order dynamic systems. The statistical properties of Green’s matching and other two-step methods are then established in Section 3. Moreover, we conduct extensive simulation studies in Section 4 to investigate the estimation behaviors of the two-step approaches and other competitive methods (Ramsay et al., 2007; Yang et al., 2021). Finally, we apply Green’s matching to equation discovery for a nonlinear pendulum in Section 5, and present conclusions and discussions in Section 6. The code, data, and proofs for this article can be found at https://github.com/Tan-jianbin/Statistical-Inference-in-General-order-Dynamic-Systems.
@@ 2.1 Local polynomial regression
With some conditions and a suitably chosen bandwidth $h_{k}$, $\widehat{D^{k}X}_{i}(\cdot)$ serves as a consistent estimator of ${D^{k}X}_{i}(\cdot)$ as $n\rightarrow\infty$ (Fan and Gijbels, 1995), although its statistical convergence rate may slow down as $k$ increases (Stone, 1982). To illustrate this phenomenon, we present several examples of local polynomial regressions in Figure 1. We observe that the estimated zero derivatives of curves accurately match their real shapes for different sample sizes $n$. However, the estimation accuracy diminishes for the higher-order derivatives. This discrepancy is likely due to the difficulties in capturing high-order effects of curves from limited noisy data.
@@ 2.2 Equation matching for dynamic systems
where $||\cdot||$ is the Euclidean norm and $w(\cdot)$ is a known function on $[0,C]$ to assign weights for different matching equations. The fitting procedure above is called equation matching, which serves as a general framework to avoid equation solving for parameter estimation in dynamic systems. We here review two important equation-matching methods: gradient matching and integral matching.
@@
We call the equation matching based on $\mathcal{T}^{i,K}_{t,\boldsymbol{\beta},\boldsymbol{\omega}}(\cdot)$ or $\mathcal{T}^{i,K-1}_{t,\boldsymbol{\beta},\boldsymbol{\omega},\alpha_{i}}(\cdot)$ as the order $K$ or (K1)(K-1) gradient matching, respectively, as both methods fit the parameters to the gradients of curves. Note that via the incorporation of integration, the order (K1)(K-1) gradient matching reduces one differential order compared to the order $K$ gradient matching. Nonetheless, these approaches both involve pre-smoothing of high-order derivatives for general-order systems, thereby imposing additional pre-smoothing errors. This is even inappropriate when the smooth high-order derivative does not exist, as shown in Section 4.1, Therefore, we require a new approach capable of eliminating all high-order derivatives in matching equations for high-order systems.
@@ 2.3 Generalized integral matching by Green’s function
In this subsection, we propose an efficient approach called Green’s matching for parameter estimation, which removes the pre-smoothed derivatives via Green’s functions of differential operators (Duffy, 2015). In general, let $\mathcal{D}$ be a differential operator of functions on $[0,C]$. We define Green’s function of $\mathcal{D}$ as a kernel function $G(t,s)$ on $0,C]^{2}$ satisfying $\mathcal{D}G(t,s)=\delta(t-s)$, where $\mathcal{D}$ is applied to the variable $t$, and $\delta(\cdot)$ is a Dirac delta function at zero. Given $G(t,s)$ of $\mathcal{D}$ and any functions $X(t)$ and $f(t)$ s.t. $\mathcal{D}X(t)=f(t)$, we have $X(t)=g(t)+\int_{0}^{C}G(t,s)f(s)\ \mathrm{d}s$, where $g(\cdot)$ is any function in the kernel space of $\mathcal{D}$, denoted as $\operatorname{Ker}(\mathcal{D}):=\left\{g(\cdot);\mathcal{D}g(t)=0,\forall t\in[0,C]\right\}$. In theory, $G(t,s)$ is considered an inverse operation for the operator $\mathcal{D}$, which would be used to reduce the differential order of the matching equations.
@@
where $\textbf{e}^{\mathbf{M}}$ is the matrix exponential of a matrix $\mathbf{M}$, $_{m,m^{\prime}}$ extracts the $(m,m^{\prime})^{th}$ element of a matrix, $\boldsymbol{\omega}_{i}=(\omega_{i0},\cdots,\omega_{i(K-1)})^{T}$, and $\boldsymbol{A}(\boldsymbol{\omega}_{i})$ is a $K\times K$ matrix with its $K^{th}$ row being $\boldsymbol{\omega}_{i}^{T}$ and $\boldsymbol{A}(\boldsymbol{\omega}_{i})]_{k,k^{\prime}}=\mathbb{I}(k=k^{\prime}-1)$ for $k\leq K-1$ and $k^{\prime}\leq K$, among $\mathbb{I}(\cdot)$ is an indicator function. Given $\boldsymbol{\psi}_{i}(t)$, $G_{i}^{K}(t,s)$ can be constructed as
@@
[…]
@@ 2.4 Optimization scheme for Green’s matching
where $\boldsymbol{Z}_{i}=(\sqrt{w(t_{1})}\hat{X}_{i}(t_{1}),\cdots,\sqrt{w(t_{H})}\hat{X}_{i}(t_{H}))^{T}$, $\boldsymbol{\Psi}_{i}=(\sqrt{w(t_{1})}\boldsymbol{a}_{i}(t_{1}),\cdots,\sqrt{w(t_{H})}\boldsymbol{a}_{i}(t_{H}))^{T}$ with $\boldsymbol{a}_{i}(t)=(\{\boldsymbol{\psi}(t)\}^{T},-\int_{0}^{C}{G}_{K}(t,s)\hat{X}_{i}(s)\mathrm{d}s,\cdots,-\int_{0}^{C}{G}_{1}(t,s)\hat{X}_{i}(s)\mathrm{d}s)^{T}$, $\boldsymbol{\nu}_{i}=(\boldsymbol{\alpha}_{i}^{T},\boldsymbol{\omega}_{i}^{T})^{T}$, and $\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})=(\sqrt{w(t_{1})}\int_{0}^{C}f_{i}(\boldsymbol{\hat{X}}(s),s;\boldsymbol{\beta})G_{K}(t_{1},s)\mathrm{d}s,\cdots,\sqrt{w(t_{H})}$ $\int_{0}^{C}f_{i}(\boldsymbol{\hat{X}}(s),s;\boldsymbol{\beta})G_{K}(t_{H},s)\allowbreak\mathrm{d}s)^{T}$. In the following, we demonstrate procedures for solving (15), which summary is given in Algorithm 1.
where $\boldsymbol{g}_{i}(\cdot)$ is a known function and $(\boldsymbol{\beta}_{1}^{T},\cdots,\boldsymbol{\beta}_{p}^{T})^{T}=\boldsymbol{\beta}$. This condition assumes that $f_{i}(\boldsymbol{X}(t),t;\boldsymbol{\beta})$ can be decomposed into a product of trajectory-related and parameter terms, a condition that is commonly adopted for statistical inference in dynamic systems (Brunton et al., 2016; Chen et al., 2017; Dai and Li, 2021). Under (16), $\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})$ in (15) satisfies $\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})=\boldsymbol{\Phi}_{i}\boldsymbol{\beta}_{i}$, where $\boldsymbol{\Phi}_{i}=(\sqrt{w(t_{1})}$ $\int_{0}^{C}\boldsymbol{g}_{i}(\boldsymbol{\hat{X}}(s),s)G_{K}(t_{1},s)\mathrm{d}s,\cdots,\sqrt{w(t_{H})}\int_{0}^{C}\boldsymbol{g}_{i}(\boldsymbol{\hat{X}}(s),s)G_{K}(t_{H},s)\allowbreak\mathrm{d}s)^{T}$. As a result, $\boldsymbol{\nu}_{i}$ and $\boldsymbol{\beta}_{i}$ can be estimated using the least squares fitting of (15) for each $i$. This approach can be applied to cases such as the gene regulatory network, oscillatory dynamic direction model, and the harmonic equations in Section 4, and also facilitates the equation discovery in Section 5.
@@
[…]
@@
For general cases, the minimizer of (15) may not have a closed form, and we instead employ iteration methods to solve the optimization problem. Letting $\boldsymbol{\nu}_{i}(\boldsymbol{\beta}):=(\boldsymbol{\Psi}_{i}^{T}\boldsymbol{\Psi}_{i})^{-1}\boldsymbol{\Psi}_{i}^{T}\allowbreak\left\{\boldsymbol{Z}_{i}-\boldsymbol{\zeta}_{i}(\boldsymbol{\beta})\right\}$, the optimization (15) can be represented as